OpenStudy (anonymous):

Show that if P ( x ) is a linear polynomial, then the composition P o P(x) is also linear and has positive slope.

terenzreignz (terenzreignz):

If P(x) is a linear polynomial, then it can be written \[\large P(x) = mx + b\]

OpenStudy (anonymous):

How would I find Find P(P(x))?

terenzreignz (terenzreignz):

Well, to find \(\large P(P(x))\) replace all instances of x in P(x) with another instance of P(x).

OpenStudy (anonymous):

So, y = m(y=mx+b) + b?

terenzreignz (terenzreignz):

y=m(mx+b)+b should suffice. Now simplify it. :)

OpenStudy (anonymous):

y=my+m^2x+mb+b?

terenzreignz (terenzreignz):

Why is there a y? \[\large P(x) = mx+b\\ \large P(P(x))=m(P(x))+b\\ \large P(P(x))=m(\color{blue}{mx+b})+b\] Please try simplifying it again.

OpenStudy (anonymous):

P=(m^2)x+bm+b

terenzreignz (terenzreignz):

Very good. \[\large P\circ P(x) = P(P(x))=\color{red}{m^2}x +bm+x\] And the slope is the coefficient of x (highlighted in red) Now, what can you say about this new slope?

terenzreignz (terenzreignz):

Sorry a slight correction \[\large P\circ P(x) = P(P(x))=\color{red}{m^2}x +bm+\color{green}b\]